# Age-of-Information-Aware Distributed Task Offloading and Resource Allocation in Mobile Edge Computing Networks

### v1

@@ Abstract
The growth in artificial intelligence (AI) technology has attracted substantial interests in age-of-information (AoI)-aware task offloading of mobile edge computing (MEC)—namely, minimizing service latency. Additionally, the use of MEC systems poses an additional problem arising from limited battery resources of MDs. This paper tackles the pressing challenge of AoI-aware distributed task offloading optimization, where user association (UA), resource allocation (RA), full-task offloading, and battery of mobile devices (MDs) are jointly considered. In existing studies, joint optimization of overall task offloading and UA is seldom considered due to the complexity of combinatorial optimization problems, and in cases where it is considered, linear objective functions such as power consumption are adopted. Revolutionizing the realm of MEC, our objective includes all major components contributing to users’ quality of experience, including AoI and energy consumption. To achieve this, we first formulate an NP-hard combinatorial problem, where the objective function comprises three elements: communication latency, computation latency, and battery usage. We derive a closed-form RA solution of the problem; next, we provide a distributed pricing-based UA solution. We simulate the proposed algorithm for various vision and language AI tasks. Our numerical results show that the proposed method Pareto-dominates baseline methods. More specifically, the results demonstrate that the proposed method can outperform baseline methods by 1.62 times smaller AoI with 41.2% less energy consumption.
@@
Age of information, mobile edge computing, resource allocation, user association, task offloading, energy efficiency, delay minimization, edge AI.
@@ 1 Introduction
AS artificial intelligence (AI) technology advances and support for AI capabilities in mobile devices (MDs) continues to grow, the demand for AI functions in edge computing is anticipated to surge. We have already experienced the impact of readily available AI applications such as voice-activated virtual assistants, generative model-based applications, and autonomous driving systems. AI applications’ high demand for computing power and energy poses a significant challenge for MDs, as they are limited in processing power and battery life. A viable solution to this problem is the implementation of task offloading using mobile edge computing (MEC), which entails shifting local tasks to a remote edge server. While this approach lightens the computational burden on MDs, it increases the communication demands of the network. For edge servers, this results in an overall increase in both computational and communication loads, leading to potential service delays. Such delays in MEC environments can degrade the quality of services such as age of information (AoI). To minimize the AoI of the data while maintaining sustainability of MEC system devices, it is essential to jointly consider communication, computation resources, and energy consumption. Addressing these concerns, especially in systems heavily reliant on AI, introduces a set of novel challenges to be tackled in optimizing task offloading.
@@
Novelty and Related Works Ours [1] [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] (’23) (’23) (’23) (’23) (’23) (’22) (’22) (’22) (’22) (’22) (’21) (’20) Architecture Multi-Cell Networks - - - - - - Multi-User Networks User Association - - - - - - - - Resource Allocation Full Offloading - - - - - - - Objective Energy Efficiency - - - - - Delay Minimization - Approach Optimization - - - - - - Lyapunov - - - - - - - - - Deep Learning - - - - - - - - - Performance Analysis Latency Minimization - - - - Energy Consumption - - - - - - Theoretical Analysis Convergence Analysis - - - Upperbound Analysis - - - - - - - - - -
1.1 Challenges for Task Offloading in MEC Systems
@@
The process of offloading local tasks in a MEC network is analogous to the conventional user association and resource allocation (UARA) problem commonly encountered in wireless networks. Typically, in a UARA scenario, a user equipment (UE) connects to a base station (BS), and resources such as bandwidth are allocated accordingly. In a MEC environment, particularly in dense or heterogeneous networks, multiple base stations are available for an MD to connect with. Similar to traditional UARA scenarios, each MD associates with an edge server (ES), receiving allocations of both computational and communication resources. However, in MEC systems, the diversity in computational capabilities of both local and edge devices, such as variations in floating point operations per second (FLOPS), VRAM capacity, and the number of available cores at the ESs, adds complexity to the MEC scenario. Additionally, in edge computing networks, delays incur from both computation and communication stages. Therefore, a joint approach to user association (UA) and resource allocation (RA) is necessary to effectively address the implications of these dual factors.
3) Trade-off between AoI and energy consumption.
@@ 1.2 Summary of Our Contributions
To the best of the authors’ knowledge, our work represents the first attempt at optimizing the joint problem of UARA for communication and computation resources in the context of distributed AI task offloading. The contributions of our work are multifaceted and can be summarized as follows:
  • We unveil a groundbreaking AI task offloading strategy, optimization UARA in a novel approach. This strategy not only i) prioritizes full-task offloading with a keen eye on the MD’s battery level but also ii) features a fully distributed optimization of load balancing.
  • We take on the challenge of the mixed-integer combinatorial optimization problem, which has not been solved in previous studies. To solve the problem, we leverage a novel auxiliary variable-based relaxation technique that enables us to balance loads by a continuous variable, not a combinatorial one. The solution we derive facilitates distributed optimization while simultaneously lowering the execution complexity.
@@
  • Through experimental results, we demonstrate that the proposed scheme consumes at least 39.1% less energy, while achieving a reduction in latency by at least 1.98 times. Best results in each objective show 5.29 times reduction in latency and 64.0% reduction in energy consumption. These findings are supported by simulations in practical environments, incorporating various AI services and load scenarios.
2 Related Works
2.1 Conventional Offloading Approaches: Pricing-Based Approach
Several studies [15, 16, 17, 18, 19, 20] have been proposed for service/load offloading in cellular networks, characterized by ultra-dense networks, massive MIMO networks, heterogeneous networks (HetNets), and cloud random access networks (C-RAN). In [21], the authors proposed a modern approach for network load balancing. The aim of the approach is distributed optimization of UARA in HetNets. In the work, the concept of pricing-based load balancing has been proposed for the first time, which enables distributed network optimization without any additional information exchange between ESs. However, such approach was limited to an objective function, specifically proportional fairness, where an artistic formulation led to a pricing framework. For instance, in [22], a pricing-based UARA optimization is further extended to backhaul-limited cellular networks.
2.2 Task Offloading in MEC
@@
TABLE II: Interpretation of the Notations Used in this Paper
@@ 3.1 System Model Overview
Figure 1 illustrates a MEC network composed of $N_{\text{ES}}$ edge servers and $N_{\text{MD}}$ MDs, where each of the edge servers has its capacity-limited computing hardware. The total computation capacities of MD $i$ and ES $j$ are denoted as $F_{\text{MD},i}$ and $F_{\text{ES},j}$ respectively. We assume that each MD always has a task at hand, creating a task whenever the previous task is finished. The generated AI service at MD $i$ has a computation load of $f_{i}$ flops and a communication load of $d_{i}$ bits. For readability, we summarize the variables used in this paper in Table II.
@@ 3.3 Communication Model
Age of Information Injected from Communication.
AoI is an important measure for many AI tasks, especially when they require real-time computation. In our scenario, AoI is incurred by 2 major components: communication latency and computation latency. Here, we examine the communication latency model. If an MD decides to process its task locally, there exists no communication delay. On the other hand, once an MD decides to offload a task to ES $j$, $d_{i}$ bits of data should be transferred to ES $j$ via a wireless channel before the service is processed.
@@ 3.4 Computing Model
Here, we introduce our computing model and provide computing delay of local computation and edge computation. In general, local devices such as mobile devices and laptops have fewer computation cores compared to the ES (GPU server). For a more practical modeling of computing delay, we divide the computation of AI tasks into parallelizable and non-parallelizable parts. Since our objective is to minimize end-to-end delay, if parallelizable portion is high, offloading can lead to higher delay reduction due to abundant number of processors at the edge server. To model this delay model, we adopt Amdahl’s law. Amdahl’s law describes the relationship between the latency of a task and the number of available processors, accounting for the portion of the task where parallel computing is feasible:
@@ 3.6 Problem Formulation
where $\alpha$ is the balance coefficient between MD energy preservation and average latency. In our problem, we do not incorporate energy conservation of ESs as an objective for minimization, given that each ES is typically connected to an energy source. Consequently, we prioritize on energy conservation of MDs with the risk of battery depletion due to limited battery capacity.
@@ 4.3 User Association
In Sections 4.1 and 4.2, we obtain the optimal solutions of $\mathbf{Y}$ and $\mathbf{Z}$ for given UA variable $\mathbf{X}$. Here, we aim to solve the optimization problem $\mathcal{P}_{1}$, with the solutions for RA variables $\mathbf{Y}$ and $\mathbf{Z}$ are provided in (22) and (24), respectively.
@@
$~{}~{}~{}~{}~{}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}\right)^{2}$ ($\mathcal{P}_{5}$a)
@@
where $c_{ij}=-D_{\text{ES}(p),ij}-D_{\text{MD},i}+\alpha\frac{G_{i}}{B_{i}}$.
@@
Before solving the problem; the challenge is that the first and second terms in Equation ($\mathcal{P}_{5}$a) are non-linear to the binary variable $x_{ij}$ and cannot be decomposed. To obtain solution, we may require computation-intensive exhaustive search, the computational complexity of which is $\mathcal{O}(N_{\text{ES}}^{N_{\text{MD}}+1}N_{\text{MD}})$. To resolve this issue, we propose a method to add two slack variables: i) $a_{j}=\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}\right)^{2}$ and ii) $b_{j}=\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)^{2}$, where $\mathbf{a}=[a_{1},\cdots,a_{N_{\text{ES}}}]$ and $\mathbf{b}=[b_{1},\cdots,b_{N_{\text{ES}}}]$. With these slack variables, we can reformulate the problem $\mathcal{P}_{5}$ without loss of generality, as follows:
@@
$\sqrt{a_{j}}=\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}$ ($\mathcal{P}_{6}$d)
$\sqrt{b_{j}}=\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}.$ ($\mathcal{P}_{6}$e)
@@ 4.3.1 Step 1: Minimize Lagrangian
$\mathcal{L}(\mathbf{X},\mathbf{a},\mathbf{b})=\sum_{j\in\mathcal{E}}(a_{j}+b_{j})+c_{ij}x_{ij}$ (25)
$\hskip 50.0pt+\sum_{j\in\mathcal{E}}\mu_{j}\left(-\sqrt{a_{j}}+\sum_{i}\sqrt{D_{\text{cm},ij}}x_{ij}\right)$
$\hskip 50.0pt+\sum_{j\in\mathcal{E}}\nu_{j}\left(-\sqrt{b_{j}}+\sum_{i}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right),$
@@
$x_{ij}^{*}=\begin{cases}1,&j=\operatorname*{argmin}(\mu_{j}\sqrt{D_{\text{cm},ij}}+\nu_{j}\sqrt{D_{\text{ES}(s),ij}}\\ &\hskip 30.0pt+c_{ij}),\\ 0,&\text{otherwise.}\end{cases}$ (29)
@@ 4.3.2 Step 2: Formulate Lagrangian dual problem
$+\sum_{i\in\mathcal{M}}\min_{j\in\mathcal{E}}(\mu_{j}\sqrt{D_{\text{cm},ij}}+\nu_{j}\sqrt{D_{\text{ES}(s),ij}}+c_{ij}).$
@@
4.3.3 Step 3: Solve Lagrangian dual problem
$\mu_{j}\leftarrow\mu_{j}+\eta_{1}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}^{*}-\frac{\mu_{j}}{2}\right)$ (31)
$\nu_{j}\leftarrow\nu_{j}+\eta_{2}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}^{*}-\frac{\nu_{j}}{2}\right),$
where $\eta_{1}$ and $\eta_{2}$ denote the step sizes for $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, respectively. As we mentioned before, the update procedure in (31) can be done with local information only; thus, the proposed method can optimize the load of the network in a distributed manner.

### v2

# Distributed Task Offloading and Resource Allocation for Latency Minimization in Mobile Edge Computing Networks

@@ Abstract
The growth in artificial intelligence (AI) technology has attracted substantial interests in latency-aware task offloading of mobile edge computing (MEC)—namely, minimizing service latency. Additionally, the use of MEC systems poses an additional problem arising from limited battery resources of MDs. This paper tackles the pressing challenge of latency-aware distributed task offloading optimization, where user association (UA), resource allocation (RA), full-task offloading, and battery of mobile devices (MDs) are jointly considered. In existing studies, joint optimization of overall task offloading and UA is seldom considered due to the complexity of combinatorial optimization problems, and in cases where it is considered, linear objective functions such as power consumption are adopted. Revolutionizing the realm of MEC, our objective includes all major components contributing to users’ quality of experience, including latency and energy consumption. To achieve this, we first formulate an NP-hard combinatorial problem, where the objective function comprises three elements: communication latency, computation latency, and battery usage. We derive a closed-form RA solution of the problem; next, we provide a distributed pricing-based UA solution. We simulate the proposed algorithm for various resource-intensive tasks. Our numerical results show that the proposed method Pareto-dominates baseline methods. More specifically, the results demonstrate that the proposed method can outperform baseline methods by 1.62 times shorter latency with 41.2% less energy consumption.
@@
Latency minimization, delay minimization, mobile edge computing, resource allocation, user association, task offloading, energy efficiency, edge AI.
@@ 1 Introduction
AS artificial intelligence (AI) technology advances and support for such technology in mobile devices (MDs) continues to grow, the demand for resource-intensive tasks in edge computing is anticipated to surge. We have already experienced the impact of emerging applications with heavy computational load such as voice-activated virtual assistants, generative model-based applications, and autonomous driving systems. These applications’ high demand for computing power and energy poses a significant challenge for MDs, as they are limited in processing power and battery life. A viable solution to this problem is the implementation of task offloading using mobile edge computing (MEC), which entails shifting local tasks to a remote edge server. While this approach lightens the computational burden on MDs, it increases the communication demands of the network. For edge servers, this results in an overall increase in both computational and communication loads, leading to potential service delays. To minimize the latency while maintaining sustainability of MEC system devices, it is essential to jointly consider communication, computation resources, and energy consumption. Addressing these concerns, especially in systems heavily reliant on AI, introduces a set of novel challenges to be tackled in optimizing task offloading.
@@
Novelty and Related Works Ours [1] [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] [13] (’23) (’23) (’23) (’23) (’23) (’23) (’22) (’22) (’22) (’22) (’22) (’21) (’20) Architecture Multi-Cell Networks Multi-User Networks User Association Resource Allocation Full Offloading Distributed optimization (multi-cell) - - - - - - Objective Energy Efficiency - Delay Minimization - Approach Optimization - Lyapunov Deep Learning - Performance Analysis Latency Minimization - Energy Consumption - Theoretical Analysis Convergence Analysis Upperbound Analysis
1.1 Challenges of Task Offloading in MEC Systems
@@
The process of offloading local tasks in a MEC network is analogous to the conventional user association and resource allocation (UARA) problem commonly encountered in wireless networks. In typical wireless networks, a user equipment (UE) connects to a base station (BS), and resources such as bandwidth are allocated accordingly. Similarly, in MEC systems, each MD associates with an edge server (ES) to receive resource allocations. However, MEC systems need to manage both computational and communication resources, where the delays arise from both computation and communication stages. What makes the problem worse is the diversity in computational capabilities of both local and edge devices, such as variations in floating point operations per second (FLOPS), VRAM capacity, and the number of available cores at the ESs. Additionally, centralized optimization, which has been adopted in most existing works, introduces further delays due to the communication overhead that occurs during information exchange. Therefore, a distributed approach to joint optimization of user association (UA) and resource allocation (RA) is necessary to effectively address the implications of these dual factors.
3) Trade-off between latency and energy consumption.
@@ 1.2 Summary of Our Contributions
To the best of the authors’ knowledge, our work represents the first attempt at optimizing the joint problem of UARA for communication and computation resources in the context of distributed task offloading. Previous works’ contributions and their limitations can be summarized as follows:
  • As in Table I, topics regarding UARA in MEC networks have been studied intensively [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Amongst them, only a few works have addressed UARA for full offloading in MEC networks [3, 7, 8, 11, 13], and even fewer research have jointly considered energy efficiency in MEC networks [3, 11].
  • More importantly, distributed optimization, which requires extremely minimal information exchange for control information, is challenging for reducing communication overhead. Despite its importance, only one work [9] in Table I has considered distributed optimization, which does not consider full-offloading.
  • One might assume that partial offloading is a more general form of full offloading, similar to how continuous variables encompass discrete variables. However, at least from an optimization standpoint, continuous-valued variables are often easier to manage because most problems can be solved using gradient-based algorithms. As a result, previous studies have tackled the full-offloading problem by relaxing it into a more tractable form with continuous-valued variables [3], which has loss of generality.
We further highlight previous studies and their contributions in Section 6. Building upon these foundations, the contributions of our work are multifaceted and can be summarized as follows:
  • We unveil a groundbreaking task offloading strategy, optimization UARA in a novel approach. This strategy not only i) prioritizes full-task offloading with a keen eye on the MD’s battery level but also ii) features a fully distributed optimization of load balancing.
  • For the sake of full-task offloading, the formulated optimiation problem is a mixed integer problem. We take on the challenge of the mixed-integer combinatorial optimization problem, which has not been solved in previous studies. To solve the problem, we leverage a novel auxiliary variable-based relaxation technique that enables us to balance loads by a continuous variable, not a combinatorial one. The solution we derive facilitates distributed optimization while simultaneously lowering the execution complexity.
@@
  • As well as theoretical analysis, our distributed optimization method is scalable for increasing network size (number of edge devices or servers), since the computational complexity at each MD and ES is $\mathcal{O}(N_{\textrm{ES}})$ and $\mathcal{O}(N_{\textrm{MD}})$, respectively. Furthermore, the total communication overhead is linearly-proportional to the number of edge servers.
  • Through experimental results, we demonstrate that the proposed scheme consumes at least 39.1% less energy, while achieving a reduction in latency by at least 1.98 times. Best results in each objective show 5.29 times reduction in latency and 64.0% reduction in energy consumption. Also, the duality gap analysis shows that the proposed method can closely achieve its latency lower bound. These findings are supported by simulations in practical environments, incorporating various AI services and load scenarios.
@@ 2.1 System Model Overview
Figure 1 illustrates a MEC network composed of $N_{\text{ES}}$ edge servers and $N_{\text{MD}}$ MDs, where each of the edge servers has its capacity-limited computing hardware. The total computation capacities of MD $i$ and ES $j$ are denoted as $F_{\text{MD},i}$ and $F_{\text{ES},j}$ respectively. We assume that each MD always has a task at hand, creating a task whenever the previous task is finished. The generated service at MD $i$ has a computation load of $f_{i}$ flops and a communication load of $d_{i}$ bits. For readability, we summarize the variables used in this paper in Table V.
@@ 2.3 Communication Model
Latency Injected from Communication.
The latency is an important measure for many AI tasks, especially when they require real-time computation. In our scenario, the Latency is incurred by 2 major components: communication delay and computation delay. Here, we examine the communication latency model. If an MD decides to process its task locally, there exists no communication delay. On the other hand, once an MD decides to offload a task to ES $j$, $d_{i}$ bits of data should be transferred to ES $j$ via a wireless channel before the service is processed.
@@ 2.4 Computing Model
Here, we introduce our computing model and provide computing delay of local computation and edge computation. In general, local devices such as mobile devices and laptops have fewer computation cores compared to the ES (GPU server). For a more practical modeling of computing delay, we divide the computation of resource-intensive tasks into parallelizable and non-parallelizable parts. Since our objective is to minimize end-to-end delay, if parallelizable portion is high, offloading can lead to higher delay reduction due to abundant number of processors at the edge server. To model this delay model, we adopt Amdahl’s law. Amdahl’s law describes the relationship between the latency of a task and the number of available processors, accounting for the portion of the task where parallel computing is feasible:
@@ 2.6 Problem Formulation
Figure 2: Illustration of the time-evolution of the AoI in our system. The communication and computation periods are depicted with purple and green slashes, respectively. The local computing is depicted with gray slashes.
Remark 3 (Objective function from the viewpoint of AoI)
In Fig. 2, we have depicted an illustration of the time-evolution of the AoI in our system. We focus on the peak AoI (PAoI) represented in [18]. The PAoI is formulated as
$\Delta=\lim_{N\rightarrow\infty}\frac{1}{N}T_{n},$ (20)
where $N$ denotes the number of processed tasks and $T_{n}$ denotes the peak of the age of $n$-th process. As depicted in Fig. 2, the peak age is proportional to the average delay of each service. Hence, with the linear age (tproportional-$\propto t$), our objective function complies with the concept of AoI with linear age and PAoI.
@@
where $\alpha$ is the balance coefficient between MD energy preservation and average latency. Specifically, the energy-related term $\frac{G_{i}}{B_{i}}$ is integrated into the time-domain objective function with the help $\alpha$. Here, $\alpha$ can be regarded as a unit conversion factor, converting energy consumption into delay. Additionally, $\alpha$ serves as a weight term of energy expenditure. By adjusting the value of $\alpha$, the network can regulate all the MDs’ local energy consumption. In our problem, we do not incorporate energy conservation of ESs as an objective for minimization, given that each ES is typically connected to an energy source. Consequently, we prioritize on energy conservation of MDs with the risk of battery depletion due to limited battery capacity.
@@ 3.3 User Association
In Sections 3.1 and 3.2, we obtain the optimal solutions of $\mathbf{Y}$ and $\mathbf{Z}$ for given UA variable $\mathbf{X}$. Here, we aim to solve the optimization problem $\mathcal{P}_{1}$, with the solutions for RA variables $\mathbf{Y}$ and $\mathbf{Z}$ are provided in (23) and (25), respectively.
@@
$~{}~{}~{}~{}~{}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)^{2}$ ($\mathcal{P}_{5}$a)
@@
where $c_{ij}=-D_{\text{ES}(p),ij}-D_{\text{MD},i}+\alpha\frac{G_{i}}{B_{i}}$, $\hat{D}_{\text{ES}(s),ij}=\frac{f_{i}\rho_{i}}{F_{\text{ES},j}Z_{\text{ES},j}}$, and $\hat{D}_{\text{cm},ij}=d_{i}/R_{ij}$
@@
Before solving the problem; the challenge is that the first and second terms in Equation ($\mathcal{P}_{5}$a) are non-linear to the binary variable $x_{ij}$ and cannot be decomposed. To obtain a solution, we may require computation-intensive exhaustive search, the computational complexity of which is $\mathcal{O}(N_{\text{ES}}^{N_{\text{MD}}+1}N_{\text{MD}})$. To resolve this issue, we propose a method to add two slack variables: i) $a_{j}=\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)^{2}$ and ii) $b_{j}=\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)^{2}$, where $\mathbf{a}=[a_{1},\cdots,a_{N_{\text{ES}}}]$ and $\mathbf{b}=[b_{1},\cdots,b_{N_{\text{ES}}}]$. With these slack variables, we can reformulate the problem $\mathcal{P}_{5}$ without loss of generality, as follows:
@@
$\sqrt{a_{j}}=\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}$ ($\mathcal{P}_{6}$d)
$\sqrt{b_{j}}=\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}.$ ($\mathcal{P}_{6}$e)
@@ 3.3.1 Step 1: Minimize Lagrangian
$\mathcal{L}(\mathbf{X},\mathbf{a},\mathbf{b})=\sum_{j\in\mathcal{E}}(a_{j}+b_{j})+\sum_{i\in\mathcal{M}}\sum_{j\in\mathcal{E}}c_{ij}x_{ij}$ (26)
$\hskip 50.0pt+\sum_{j\in\mathcal{E}}\mu_{j}\left(-\sqrt{a_{j}}+\sum_{i}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)$
$\hskip 50.0pt+\sum_{j\in\mathcal{E}}\nu_{j}\left(-\sqrt{b_{j}}+\sum_{i}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right),$
@@
$x_{ij}^{*}=\begin{cases}1,&j=\operatorname*{argmin}(\mu_{j}\sqrt{\hat{D}_{\text{cm},ij}}+\nu_{j}\sqrt{\hat{D}_{\text{ES}(s),ij}}\\ &\hskip 30.0pt+c_{ij}),\\ 0,&\text{otherwise.}\end{cases}$ (30)
@@ 3.3.2 Step 2: Formulate Lagrangian dual problem
$+\sum_{i\in\mathcal{M}}\min_{j\in\mathcal{E}}(\mu_{j}\sqrt{\hat{D}_{\text{cm},ij}}+\nu_{j}\sqrt{\hat{D}_{\text{ES}(s),ij}}+c_{ij}).$
@@
3.3.3 Step 3: Solve Lagrangian dual problem
$\mu_{j}\leftarrow\mu_{j}+\eta_{1}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}^{*}-\frac{\mu_{j}}{2}\right)$ (32)
$\nu_{j}\leftarrow\nu_{j}+\eta_{2}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}^{*}-\frac{\nu_{j}}{2}\right),$
where $\eta_{1}$ and $\eta_{2}$ denote the step sizes for $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, respectively. As previously stated, the update procedure in (32) can be done with local information only; thus, the proposed method can optimize the load of the network in a distributed manner.
Now, we analyze the information exchange of the network. The information exchange necessary for the proposed method can be outlined as follows:
  • Control information collection from the MDs to the ESs: From the update equation (32), the information required for ES $j$ to determine its price is $\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\textrm{cm},ij}}x_{ij}^{*}$ and $\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\textrm{ES}(s),ij}}x_{ij}^{*}$. Those values are already accessible at each ES, as ES $j$ can obtain the delay term of the serving users ($x_{ij}^{*}=1$) without additional information exchange.
  • ED-side actions: The task of each MD is to simply compare the weighted delay for ESs and determine their $x_{ij}$ values, as outlined in (30). Subsequently, the remaining communication load is negligible since the MD only needs to send two integers when the MEC network is initialized – each representing the number of bits and the number of flops of the task.
  • Information exchange between the ESs: No information exchange between the ESs is needed, as the price of an ES is calculated based on the load from the allocated MDs only.
  • Price broadcasting from the ESs to the MDs: As the price each ES needs to broadcast is a single floating point number, only a short communication is necessary.